# Word Embeddings Revisited: Do LLMs Offer Something New?

### v1

@@ 1 Introduction
While LLMs have shown remarkable success across various well-known NLP tasks Bubeck et al. (2023); Dai et al. (2022); Du et al. (2022); Smith et al. (2022), it remains an open question if their enhanced performance can solely be attributed to their larger scale or if the embeddings they generate are fundamentally distinct from those created by traditional encoding models such as Sentence-BERT or Universal Sentence Encoders. Since the introduction of the Transformer architecture, model development has exploded with innovations, improving year after year. However, the medium has remained the same: written language. In theory, the similarity and relatedness of words have not changed, but the way that models treat words and their similarity has. More recently, Fournier et al. (2020) evaluated smaller transformer models like GPT-2 and BERT on analogy tasks and found that these models performed poorly on Encyclopedic and Lexicographic analogies, further raising the question of how embedding performance changes as model size increases. As such, we perform the following two analyses in this paper.
@@ 2.1 Embeddings
A sampling of models from the last five years was used to create a multitude of embeddings for experimentation. We classify these models into two broad categories: Large Language Model (LLM) based – models with over 1B parameters and “Classical” models – those with under 1B parameters. Tested here are LLaMA2-7B (dim $=4096$) from Meta AI Touvron et al. (2023), OpenAI’s embedding model ADA-002 (dim $=1536$), and Google’s PaLM2 embedding model Gecko-001 (dim $=768$Anil et al. (2023). To more clearly see the differences between these models and older (“classical”) ones, Meta AI’s LASER (dim $=1024$Artetxe and Schwenk (2019), Universal Sentence Encoder (USE) (dim $=512$Cer et al. (2018), and SentenceBERT (SBERT) (dim $=384$Reimers and Gurevych (2019) were also tested.
@@ 2.5 Do LLMs Offer Something New?
In order to determine whether LLMs offer something new/very different, we investigated how embedding models (LLMs and Classical) agreed or disagreed with each other regarding the similarity of word pairs. To be more specific, we computed the difference in rank between two related words for each pair of embedding models. For example, if the 5th closest word to “bad” according to ADA-002’s embedding was “worst”, while "worst" was the 10th nearest word to “bad” according to LLaMA’s embeddings, we would compute a difference of $-5$ for that word pair while comparing ADA vs. LLaMA. If two models tend to agree on the relative similarity of word pairs most of the time, we would expect a mean of 00 for this value and a small variance.
@@ Appendix A Appendix
Model Analogy Method 1. Inflectional Morphology 2. Derivational Morphology 3. Encyclopedic Semantics 4. Lexicographic Semantics

### v2

@@ 1 Introduction
While LLMs have shown remarkable success across various well-known NLP tasks Bubeck et al. (2023); Dai et al. (2022); Du et al. (2022); Smith et al. (2022); Sarkar et al. (2023); Akter et al. (2023), it remains an open question if their enhanced performance can solely be attributed to their larger scale or if the embeddings they generate are fundamentally distinct from those created by traditional encoding models such as Sentence-BERT or Universal Sentence Encoders Mahajan et al. (2023); Sarkar et al. (2022). Since the introduction of the Transformer architecture, model development has exploded with innovations, improving year after year. However, the medium has remained the same: written language. In theory, the similarity and relatedness of words have not changed, but the way that models treat words and their similarity has. More recently, Fournier et al. (2020) evaluated smaller transformer models like GPT-2 and BERT on analogy tasks and found that these models performed poorly on Encyclopedic and Lexicographic analogies, further raising the question of how embedding performance changes as model size increases. As such, we perform the following two analyses in this paper.
@@ 2.1 Embeddings
A sampling of models from the last five years was used to create a multitude of embeddings for experimentation. We classify these models into two broad categories: Large Language Model (LLM) based – models with over 1B parameters and “Classical” models – those with under 1B parameters. Tested here are LLaMA2-7B (dim =4096=4096) from Meta AI Touvron et al. (2023), OpenAI’s embedding model ADA-002 (dim =1536=1536), and Google’s PaLM2 embedding model Gecko-001 (dim =768=768Anil et al. (2023). To more clearly see the differences between these models and older (“classical”) ones, Meta AI’s LASER (dim =1024=1024Artetxe and Schwenk (2019), Universal Sentence Encoder (USE) (dim =512=512Cer et al. (2018), and SentenceBERT (SBERT) (dim =384=384Reimers and Gurevych (2019) were also tested.
@@ 2.5 Do LLMs Offer Something New?
In order to determine whether LLMs offer something new/very different, we investigated how embedding models (LLMs and Classical) agreed or disagreed with each other regarding the similarity of word pairs. To be more specific, we computed the difference in rank between two related words for each pair of embedding models. For example, if the 5th closest word to “bad” according to ADA-002’s embedding was “worst”, while "worst" was the 10th nearest word to “bad” according to LLaMA’s embeddings, we would compute a difference of 5-5 for that word pair while comparing ADA vs. LLaMA. If two models tend to agree on the relative similarity of word pairs most of the time, we would expect a mean of 0 for this value and a small variance.
@@ Appendix A Appendix
Model Analogy Method
1. Inflectional
Morphology
2. Derivational
Morphology
3. Encyclopedic
Semantics
4. Lexicographic
Semantics